Two 1.5 kg masses are 1.5 m apart (center to center) on a frictionless table. Each has + 9.1 μC of charge.

What is the magnitude of the electric force on one of the masses?

What is the initial acceleration of the mass if it is released and allowed to move?
Answer #1

Solution)

We know,

a) F = k q1q2 / r^2

= 9 x 10^9 * (9.1 x 10^-6)^2 / 1.5^2

= 0.33124 N

======

b) a = F / m = 0.33124 / 1.5 = 0.2208 m/s^2

======